OpenStudy (anonymous):

HELP PLEASE!!! WILL GIVE A MEDAL!! Given the linear equation 4y = 8x – 3, find the y-intercept. A.) -3/4 B.) -2 C.) 3/4 D.) 2

OpenStudy (anonymous):

@beautifulll @GotToLoveJenny @TechnoSoul

OpenStudy (anonymous):

b=-3/4

OpenStudy (anonymous):

B. -3/4

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i pick b=-3/4

OpenStudy (anonymous):

y=mx+b, b is your y intercept, because you have a value in front of y, you have to divide the rest of the equation by that value, so you'd get y=2x-3/4. -3/4 is your answer

OpenStudy (anonymous):

A i mean

OpenStudy (anonymous):

thanks SO much!! u guys r lifesavers!! lolz

OpenStudy (anonymous):

It would not be A because, y=mx+b. the + b is what's important here. because of that, even if your answer is negative, it will still stay negative.

OpenStudy (anonymous):

k thank